OpenStudy (anonymous):

Find the degree of the polynomial function. f(x) = -2x + 7x^2

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Can someone help me?

OpenStudy (anonymous):

Find the degree of the polynomial function. f(x) = -2x + 7x^2 help

OpenStudy (anonymous):

help help help help help help

OpenStudy (anonymous):

Do anyone know how to find the degree of the polynomial?

OpenStudy (anonymous):

f(x) = -2x + 7x^2 when this equation is put into standard form it will be f(x) = 7x^2 -2x The degree is the exponent of the first term

OpenStudy (anonymous):

the first term is 7x^2 therefore the degree would be 2 (This polynomial would be a quadratic binomial because it has a degree of two and has two terms)

OpenStudy (anonymous):

so the degree of the polynomial function would be 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the degree of the polynomial function. f(x) = πx5 - 6x4 - 9

OpenStudy (anonymous):

Can you help me with this @Little-Ho

OpenStudy (anonymous):

f(x) = πx^5 - 6x^4 - 9

OpenStudy (anonymous):

sorry just got back

OpenStudy (anonymous):

ok so the equation that you now have is great because it is already in standard form f(x) = πx^5 - 6x^4 - 9 πx^5 is the first term with the highest exponent that exponent (5) is the degree of this polynomial function